OpenStudy (anonymous):

Evaluate -s^4 when s= -3

OpenStudy (lgbasallote):

substiute s=-3 \[\implies -(-3)^4\] so do you know the value of (-3)^4?

OpenStudy (anonymous):

My issue is that I'm not sure if the problem will translate into -3^4 or 3^4

OpenStudy (lgbasallote):

it translates to -3^4

OpenStudy (lgbasallote):

because the negative sign is outside of the parenthesis..make sense?

OpenStudy (anonymous):

Even if the original problem does not contain parenthesis?

OpenStudy (lgbasallote):

hmm...in the original problem what you're raising to 4 is s right? and it is given that s = -3 so you're going to raise -3 to the 4th power which translates to (-3)^4 because -3 is the WHOLE s...that's why you put a parenthesis. then there's a negative sign outside -(-3)^4 right?

OpenStudy (anonymous):

Well when I'm substituting -s with -3, it seems to me that those two negative signs would cancel each other out?

OpenStudy (anonymous):

Sorry it took me so long to respond- for some reason it wouldn't allow me to post!

OpenStudy (lgbasallote):

noooo

OpenStudy (lgbasallote):

remember PEMDAS

OpenStudy (lgbasallote):

you raise s = -3 to 4 first THEN multiply to -1

OpenStudy (lgbasallote):

would you like to see an example?

OpenStudy (anonymous):

Yes please!

OpenStudy (lgbasallote):

\[\huge -x^2y^3\] given that x= -4 and y = 2 so if you substitute.. \[\huge \implies -(-4)^2 (2)^3\] remember that: \[\huge (-4)^2 \implies -4 \times -4 \implies 16\] and \[\huge (2)^3 \implies 2\times 2 \times 2 \implies 8\] so... \[\huge -(-4)^2 (2)^3 \implies -(16)(8) \implies -128\]

OpenStudy (lgbasallote):

did that help?

OpenStudy (anonymous):

YES!!!! OK I see it now!!! Thank you!

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

So it would be -(-3)^4 which would be -(81) and my final answer would be -81

OpenStudy (lgbasallote):

right!!!

OpenStudy (anonymous):

Awesome! Thanks again!

OpenStudy (lgbasallote):

welcome